OpenStudy (mwhit2101):

How to graph this equation? y=1/2x+2 So I obviously know how to go to desmos.com and type in that equation and get an answer. But I can't do that during exams. I need to knoew how to literally put this on the graph. And I don't. So can someone please go through the steps with me and show me how to?

OpenStudy (kamille):

is it \[y=\frac{ 1 }{ 2x+2}\] or \[y=\frac{ 1 }{ 2 }x+2\]

OpenStudy (mwhit2101):

Second one

OpenStudy (kamille):

well, it is linear function, right? Linear function graph is a line, so all you need is two points. To find those points (so you can make a graph), you need to take any number and insert it in the place of x - that way you will find y. For example, if you take 0: y=1/2*0+2=2 so your point is (0;2) another x you take can be 1, so you get y=1/2*1+2=2.5 so another point is (1;2.5) So now you have to points A (0;2) and B (1;2.5). Put those points on the graph and draw the line through these points